OpenStudy (anonymous):

Find the maximum or minimum of the following quadratic function: y =x2 - 10x + 25

OpenStudy (jellisoncollege):

set y to zero 0=x^2-10x+25 0=(x-5)(x-5) 0=x-5 (move 5 to the other side) x=5 Maximum and minimum is 5.

OpenStudy (anonymous):

thanks !